OpenStudy (anonymous):

how do i factor these please help 3x3 + 9x2 + x + 3?

OpenStudy (anonymous):

3x3 = 9 + 9x2 = 18 = x = 18x + 3 = 21x start from there

OpenStudy (bradely):

3x^3 + 9x2 + x + 3 3x^2(x+3)+1(x+3) (x+3)(3x^2+1) Source: http://www.mathskey.com/question2answer/ask

OpenStudy (anonymous):

thank you :)